1.      The prices for LED are listed: 22 500, 24 000, 21 500, 20 200, 20 600, 21 100, 210 000, 19 300, 25 000 and 22 500. Estimate the true mean of the data set with 90% confidence level.


1
Expert's answer
2022-04-05T06:46:11-0400

"m=(22500+24000+21500+20200+20600+21100+210000+19300+25000+22500)\/10=40670"

"\sigma=\sqrt{\frac{\sum (x-m)^2}{n-1}}=\sqrt{\frac{330148900+377888900+367488900+419020900+402804900+382984900+2.867x10^{10}+456676900+245548900+330148900}{9}}=59833"

"CI=m \plusmn z \frac{\sigma}{\sqrt{n}}=40670 \plusmn 1.645 \frac{59833}{\sqrt{10}}=40670 \plusmn 31124.8"